The area of a rectangular red sticker is 98 square millimeters. The perimeter is 42 millimeters. What are the dimensions of the sticker?

Answer by Boreal(15235)

A=L*W=98
Therefore L=98/W.
Perimeter is 42 mm and half the perimeter is L+W=21 mm
But L=98/W
Therefore (98/W)+W=21 mm
multiply everything by W
98+w^2=21W
w^2-21w+98=0
(w-14)(w-7)=0
w=14 or 7
L=7 or 14
By convention, L is longer, so the dimensions are 14 mm x 7 mm.